A man mixes peanuts selling at 40 cents a pound with cashew nuts selling at 70 cents a pound to make a 50-pound mixture which he sells at 52 cents a pound. How many pounds of each kind of nuts must he mix?(Only an algebraic solution will be accepted.)
Answer by nerdybill(7384)

A man mixes peanuts selling at 40 cents a pound with cashew nuts selling at 70 cents a pound to make a 50-pound mixture which he sells at 52 cents a pound. How many pounds of each kind of nuts must he mix?
.
Let x = pounds of peanuts
then
50-x = pounds of cashews
.
.40x + .70(50-x) = .52(50)
.40x + 35 -.70x = 26
35 -.30x = 26
35 = 26 + .30x
9 = .30x
30 pounds = x (peanuts)
.
Cashews:
50-x = 50-30 = 20 pounds (cashews)